OpenStudy (anonymous):

4x+3y=2 3x-4y=14 please fast

OpenStudy (kc_kennylau):

Multiply (1) by 4, (2) by 3.

OpenStudy (anonymous):

multiply the first equation with 3 and the second one with 4 3.(4x+3y=2) and 4.(3x-4y=14) subtracting second equation from first 12x+9y-12x+16y=2-14 y=-12/25 now put this value of y in any of the two equations to get the value of x

OpenStudy (anonymous):

4x+3y=2) and 4.(3x-4y=14) subtracting second equation from first 12x+9y-12x+16y=2-14 y=-12/25 now put this value of y in any of the two equations to get the value of x

OpenStudy (anonymous):

but i am not getting the correct answer

OpenStudy (anonymous):